# Over-the-Air Fusion of Sparse Spatial Features for Integrated Sensing and Edge AI over Broadband Channels

### v1

@@ III-B Radio Resource Allocation
Let $|p_{k,m}|^{2}$ denote the transmit power invested to subcarrier $m$ by agent $k$. Then, $\{|p_{k,m}|^{2}\}$ depend on the sparsity of paired voxels channel gains, and receive SNRs (after aggregation). To be specific, the agents containing zero voxel feature vectors do not participate in transmission on the designated subcarrier, i.e., $p_{k,m}=0$ if $\sum_{v=1}^{V}S_{k,v}A_{v,m}=0$. All the agents participating in the transmission on the designated subcarrier shall set transmit power to align their signal magnitude as required for AirComp[8]. It follows that the receive SNR, denoted as $\gamma_{v}$ for voxel $v$, is given as
@@ III-C Over-the-Air Fusion
where the normalization parameters $\sigma$ and $\mu$ in (10) are shared by all agents and set such that the distribution of pre-processed features is zero-mean and unit-variance. The extension to other fusions functions (e.g., max-pooling [3]) is straightforward by applying additional post- and/or pre-processing functions. The pairing matrix $\mathbf{A}$ maps the pre-processed features, $\{\tilde{x}_{k,v}[\ell]\}_{v=1}^{V}$ to the $\ell$-th symbol of each subcarrier. Then the symbol transmitted by agent $k$ over subcarrier $m$ can be written as
@@ VI Optimal VoCa-PPA: Compact Tree Design
The greedy VoCa-PPA algorithm in the preceding section is computation-efficient but sub-optimal. In this and the next sections, we present an optimal and efficient approach for solving the VoCa-PPA Problem in P1 or equivalently Problem P2. The tree-search based approach consists of two components – compact tree design in this section and fast tree search in the next section. In general, Problem P2 can be viewed as a special case of the max-linear assignment problem, and its optimal solution can be searched for using the well-known ranking method (see, e.g., [35]). The novelty of our design, which yields a higher efficiency than the existing method, lies in exploiting the special structure of Problem P2. In particular, a derived useful property of its objective leads to a dramatic reduction of the dimensionality of the search space. As a common practice in solving bipartite matching problems, assume equal numbers of voxels and subcarriers $M=V$ in the sequel without loss of generality as the case of $M>V$ can be augmented with $(M-V)$ dummy voxels with all-zero sparsity indicators for all agents.
@@ VIII-A Experimental Settings
  • OPV2V dataset. The OPV2V dataset [33] considers a vehicle-to-vehicle communication scenario where an ego vehicle fuses sensory features from helping vehicles detect other vehicles in a traffic scene. A data frame involves two to five vehicles, one of which is selected as the ego vehicle. Each vehicle captures a LiDAR point cloud of the surrounding environment and objects, which is projected onto the ego vehicle’s coordinates and processed by a PointPillar backbone into a two-dimensional local spatial feature map with $V_{\mathsf{h}}=256$ and $V_{\mathsf{w}}=352$ being the number of voxels along the height and width of the perception region, respectively. Each voxel is associated with a feature vector with dimension $L=128$, and thus the size of the local feature map is $128\times 256\times 352$. Therein, $50591$ out of $V_{\mathsf{h}}V_{\mathsf{w}}=90112$ voxels are empty over all samples in the dataset, regarded as dummy, and waived of transmission for all evaluated methods. Following[36], the ego vehicle wirelessly aggregates the feature map from all other vehicles and inputs the fused feature map into a region proposal network to obtain the vehicle detection result. The detection performance is measured by the average precision (AP) at an Intersection over Union (IoU) threshold of 0.7. It is defined as the area under the precision-recall curve resulting from the said detection model, where a detected bounding box is considered true-positive if it overlaps with a ground-truth bounding box with an IoU higher than $0.7$[11].
@@ IX Concluding Remarks
This work opens up several research directions on task-oriented communication schemes for ISEA. For example, digital Spatial AirFusion can be developed for better compatibility with existing digital systems, enabling incorporation of digital transmission techniques such as modulation and coding schemes. Another interesting topic is the interplay between Spatial AirFusion and more sophisticated physical layer techniques such as MIMO. In addition, integrating Spatial AirFusion with semantic data sourcing, which broadcasts low-dimensional queries to trigger transmission on semantically relevant agents, can further reduce communication cost[38].

### v2

@@ III-B Radio Resource Allocation
Let $|p_{k,m}|^{2}$ denote the transmit power invested to subcarrier $m$ by agent $k$. Then, $\{|p_{k,m}|^{2}\}$ depend on the sparsity of paired voxels channel gains, and receive SNRs (after aggregation). To be specific, agent $k$ does not participate in the aggregation of voxel $v$ if $S_{k,v}=0$, thereby setting its transmitting power to zero on the subcarrier designated for voxel $v$. This is mathematically given by: $p_{k,m}=0$ if $\sum_{v=1}^{V}S_{k,v}A_{v,m}=0$. All the agents participating in the transmission on the designated subcarrier shall set transmit power to align their signal magnitude as required for AirComp[8]. It follows that the receive SNR, denoted as $\gamma_{v}$ for voxel $v$, is given as
@@ III-C Over-the-Air Fusion
where the normalization parameters $\sigma$ and $\mu$ in (10) are shared by all agents and set such that the distribution of pre-processed features is zero-mean and unit-variance. The extension to other fusions functions (e.g., max-pooling [3]) is straightforward by applying additional post- and/or pre-processing functions. The pairing matrix $\mathbf{A}$ maps the pre-processed features, $\{\tilde{x}_{k,v}[\ell]\}_{v=1}^{V}$ to the $\ell$-th symbol of each subcarrier. Then the symbol transmitted by agent $k$ over subcarrier $m$ can be written as [^2] For notational simplicity, we have assumed that the features are modulated onto the real, or in-phase, component of the transmitted symbols, which is also common in many AirComp literature, e.g., [30, 27] In practice, it is possible to modulate features on both the in-phase and quadrature components, which reduces the AirComp latency by half (see, e.g., [44, 45]). The extension of our work to this case is straightforward without changing the design of our framework and algorithms.
@@
Remark 2.
(System Scalability) One key advantage of Spatial AirFusion against digital orthogonal access is its high scalability w.r.t. the number of agents and data volume. Aligned with real-world challenges, increasing the number of agents in cooperative perception is a trend in relevant literature, e.g., from $2$ agents in F-Cooper[46] to $6$ agents in OPV2V[34] to $12$ agents in V2X-Sim[47], which results in growing communication overhead for digital orthogonal access. In contrast, the increase in number of agents does not add to latency or bandwidth consumption in AirFusion thanks to simultaneous access, but also mitigates both channel and data noise as found in [2]. The increase in data volume is due to sensor advancements, e.g., LiDAR sensing resolution and range. Spatial AirFusion copes with this challenge by 1) fusion of resolution-invariant voxel features instead of raw data; 2) fusion on sparse non-empty voxels instead of all voxels in the sensing region.
@@ V-B Greedy VoCa-PPA
V-C Complexity Analysis
The time complexity of Algorithm 1 is presented as follows. Before starting the iteration, we sort $\{c_{1,m},\ldots,c_{K,m}\}$ in descending order for each $m=1,\ldots,M$ and store the results. The complexity of this step is $\mathcal{O}(MK\log K)$. Then, in the $v$-th iteration, (17) shall be evaluated. The inner $\max\limits_{k}\ {c_{k,m}}{S_{k,v}}$ is obtained with $\mathcal{O}(1)$ given $\{c_{1,m},\ldots,c_{K,m}\}$ sorted and $S_{k,v}$ binary, and the outer operation costs $\mathcal{O}(M)$. Hence, the complexity of the iterating process is $\mathcal{O}(MV)$. The total complexity of Algorithm 1 is thus $\mathcal{O}(M\max\{K\log K,V\})$.
This complexity is comparable to basic algorithms in OFDM such as Fast Fourier Transform (FFT). Moreover, our algorithm does not involve floating-point multiplications, making it highly efficient for implementations in standard hardware[48].
@@ VI Optimal VoCa-PPA: Compact Tree Design
The greedy VoCa-PPA algorithm in the preceding section is computation-efficient but sub-optimal. In this and the next sections, we present an optimal and efficient approach for solving the VoCa-PPA Problem in P1 or equivalently Problem P2. The tree-search based approach consists of two components – compact tree design in this section and fast tree search in the next section. In general, Problem P2 can be viewed as a special case of the max-linear assignment problem, and its optimal solution can be searched for using the well-known ranking method (see, e.g., [49]). The novelty of our design, which yields a higher efficiency than the existing method, lies in exploiting the special structure of Problem P2. In particular, a derived useful property of its objective leads to a dramatic reduction of the dimensionality of the search space. The motivation of organizing the search into a search tree is to reduce the search complexity by node pruning with a branch-and-bound-inspired method. The method hinges on proper selection of the branching variable and bounding the global objective with local ones, as will be introduced shortly. As a common practice in solving bipartite matching problems, assume equal numbers of voxels and subcarriers $M=V$ in the sequel without loss of generality as the case of $M>V$ can be augmented with $(M-V)$ dummy voxels with all-zero sparsity indicators for all agents.
@@ VII-C Fast Tree-Search Algorithm
VII-D Complexity Analysis
The computation complexity of visiting each node by Algorithm 2 in the defined tree can be divided into that of 1) local objective evaluation by (20), which is $\mathcal{O}(M)$; 2) pruning/stopping determination by Proposition 1 and Proposition 2, which is $\mathcal{O}(1)$; 3) enumeration of child nodes in ascending order of local objectives, which in the worst case $\mathcal{O}(N_{\text{child}}\log N_{\text{child}})$, where $N_{\text{child}}$ is the number of child nodes. Note that the last term is amortized by all $N_{\text{child}}$ nodes, and thus the amortized complexity per node is in fact upper bounded as $\mathcal{O}(\log N_{\text{child}})<\mathcal{O}(\log M!)=\mathcal{O}(M\log M)$. Meanwhile, the total number of visited nodes is upper bounded by $KN_{\text{sol}}$, where $N_{\text{sol}}$ is the number of solutions (leaf nodes) enumerated before the algorithm stops. Therefore, the worst case complexity of Algorithm 2 is $\mathcal{O}(N_{\text{sol}}KM\log M)$. In the worst case, $N_{\text{sol}}$ can still reach the size of the full solution space, which is exponential in $M$. This is inevitable due to the NP-hardness of Problem P2. However, the empirical number of solutions visited is usually substantially lower than the worst case, thanks to the proposed fast tree-search algorithms. To illustrate the empirical complexity, $95$-th percentiles of $N_{\text{sol}}$ under different $M$ and $K=4$ are presented in Table 1 along with the solution space size, i.e., the number of all possible solutions.
TABLE I: Numbers of Enumerated Solutions and All Solutions
$M$ 95-th percentile of $N_{\text{sol}}$ Number of all solutions
$8$ $12$ $1.10\times 10^{4}$
$16$ $99$ $1.20\times 10^{11}$
$32$ $653$ $1.78\times 10^{27}$
@@ VIII-A Experimental Settings
  • OPV2V dataset. The OPV2V dataset [34] considers a vehicle-to-vehicle communication scenario where an ego vehicle fuses sensory features from helping vehicles detect other vehicles in a traffic scene. A data frame involves two to five vehicles, one of which is selected as the ego vehicle. Each vehicle captures a LiDAR point cloud of the surrounding environment and objects, which is projected onto the ego vehicle’s coordinates and processed by a PointPillar backbone into a two-dimensional local spatial feature map with $V_{\mathsf{h}}=256$ and $V_{\mathsf{w}}=352$ being the number of voxels along the height and width of the perception region, respectively. Each voxel is associated with a feature vector with dimension $L=128$, and thus the size of the local feature map is $128\times 256\times 352$. We find that in all voxels observed by all agents, over 90% are empty, resulting in zero feature vectors, which conforms to the observations by [11, 31]. Therein, $50591$ out of $V_{\mathsf{h}}V_{\mathsf{w}}=90112$ voxels are empty over all samples and all agents in the dataset, regarded as dummy, and waived of transmission for all evaluated methods. The ego vehicle wirelessly aggregates the feature map from all other vehicles and inputs the fused feature map into an RPN, as in [46], to obtain the vehicle detection result. The detection performance is evaluated by comparing the downstream network output with the ground truth, measured by the average precision (AP) at an Intersection over Union (IoU) threshold of 0.7. It is defined as the area under the precision-recall curve resulting from the said detection model, where a detected bounding box is considered true-positive if it overlaps with a ground-truth bounding box with an IoU higher than $0.7$[11].
@@ IX Concluding Remarks
We acknowledge that several assumptions have been made in this paper to simplify the exposition, which motivate further studies. Identical feature variance is assumed across all agents, the relaxation of which requires feature statistics-aware power control. Symbol-level synchronization and sufficient channel coherence time across all agents are assumed to facilitate AirComp design. However, under high mobility, such conditions may not hold, requiring advanced scheduling and air-interface designs.
This work opens up several research directions on task-oriented communication schemes for ISEA. For example, digital Spatial AirFusion can be developed for better compatibility with existing digital systems, enabling incorporation of digital transmission techniques such as modulation and coding schemes. Another interesting topic is the interplay between Spatial AirFusion and more sophisticated physical layer techniques such as MIMO. In existence of strong resource heterogeneity across agents, asynchronous feature aggregation could be necessitated, and the relevant scheduling and fusion schemes warrant future studies. In addition, integrating Spatial AirFusion with semantic data sourcing, which broadcasts low-dimensional queries to trigger transmission on semantically relevant agents, can further reduce communication cost[51].